A one-year U.S. Treasury security has a nominal interest rate of 2.25 percent. If the expected real rate of interest is 1.5 percent, what is the expected annual inflation rate?
Answer #1

(1+real rate)=(1+nominal rate)/(1+inflation rate)

(1+0.015)=(1+0.0225)/(1+inflation rate)

1+inflation rate=(1.0225/1.015)

inflation rate=(1.0225/1.015)-1

=0.74%(Approx).